OpenStudy (anonymous):

Am I right? What is the vertex of the graph of y = 2|x – 3| + 5? My answer is (0,-1).

OpenStudy (anonymous):

l;;;;;; 44444444

OpenStudy (anonymous):

wut.

OpenStudy (jdoe0001):

ahemm

OpenStudy (jdoe0001):

@TheHero still there?

OpenStudy (anonymous):

Use the definitions of |x| to express the funtion as two ecuations: |x| ={ x if x≥0 {-x if x<0 So in this case 2|x – 3| + 5={ 2x-6+5 = 2x-1 if x≥3 { 6-2x+5 = 11-2x If x<3 So you have the ecuations of two lines with opposite slopes (2x-1 has positive slope and 11-2x has negative slope). So the vertex must be the intersection between the line.An we can get the x corrdinate of the vertex by solving the ecuation: 2x-1 = 11-2x

OpenStudy (anonymous):

Does that mean I'm right?

OpenStudy (jdoe0001):

lol

OpenStudy (anonymous):

That mean that the vertex is the pont (3,5)

OpenStudy (jdoe0001):

@TheHero what's the expression inside the absolute value?

OpenStudy (anonymous):

Oh.... Well, thanks ;~;

OpenStudy (anonymous):

Welcome